# Limits of Generalization in RLVR: Two Case Studies in Mathematical Reasoning

### v1

@@ 1 Introduction
Large language models (LLMs) have recently advanced rapidly on mathematical and programming benchmarks [5, 6, 21, 2, 22]. A key driver is Reinforcement Learning with Verifiable Rewards (RLVR), which fine-tunes pretrained models against automatically checkable signals such as exact answers or unit tests [5, 23]. This paradigm eliminates reliance on human annotation, enabling scalable training on large problem sets and delivering consistent gains on challenging reasoning tasks [4].
Despite strong empirical gains, the nature of RLVR’s improvements remains unclear. Studies show it often boosts accuracy while reducing exploration [20, 3, 24, 14], with base model ability acting as a ceiling [24, 20]. Gains largely reflect re-weighting existing solutions, and reasoning coverage can even contract at larger sample sizes. Standard $Pass@K$ further over-credits “lucky” completions [17], whereas stricter metrics like CoTCoT-$Pass@K$ are more reliable but harder to scale. Reward design can also yield surprising effects: a single example can rival large-scale training [16], and even spurious rewards can drive improvements in models with strong procedural biases [8]. Overall, RLVR appears to stabilize existing competencies rather than induce new reasoning strategies.
@@ 2.1 Tasks
Activity Scheduling. Each activity $i$ has start and finish times $(s_{i},f_{i})$ with $s_{i}<f_{i}$, and intervals are half-open $s_{i},f_{i})$. The goal is to select a maximum subset of non-overlapping activities [18]. Instances are constructed so that the greedy earliest-finish-time algorithm with deterministic tie-breaking yields the unique optimum, reported as IDs sorted by $f_{i}$ (ties by smaller $i$).
Longest Increasing Subsequence (LIS). Given $a_{1},\dots,a_{n}\in\mathbb{Z}$, find indices $1\leq i_{1}<\cdots<i_{k}\leq n$ maximizing $k$ with $a_{i_{1}}<\cdots<a_{i_{k}}$. Uniqueness is enforced via an $O(n^{2})$ dynamic-programming count, and the LIS is reconstructed with patience sorting in $O(n\log n)$ [19].
@@ 2.2 Reward Functions
$r_{\text{ids,exact}}(y)\;=\;\mathbb{I}\!\bigl[\hat{s}(y)=s^{\star}\bigr].$
@@
$r_{\text{ids,prefix}}(y)\;=\;\max\Bigl\{\,0,\ \frac{m(y)}{L}\;-\;\gamma\,\mathbb{I}\!\bigl[\,\hat{s}(y)=\varnothing\ \lor\ |\hat{s}(y)|\neq L\,\bigr]\Bigr\}.$
@@ 2.3 Training & Evaluation
We fine-tune Qwen2.5-7B-Instruct [12] with GRPO [9, 5] using the verl framework [11]. Unless noted, max generation length is $T_{\max}=2048$ (extended to $7680$ for LIS with $r_{\text{ans+fmt}}$). Each PPO update uses $256$ prompts with $8$ rollouts from vLLM [7], trained for $20$ epochs ($120$ updates) at learning rate $10^{-6}$ and no KL penalty ($\beta_{\text{KL}}=0$). Training prompts match evaluation format.
@@
  • Pass@k [24]. Success under a $k$-sample budget, defined by whether at least one of the $k$ generations is correct:
@@
  • Self-consistency (SC) [15]. Agreement between the majority prediction aggregated over $k$ generations and the ground truth:
@@ 3.1 Training with Exact Answer Reward
Activity Scheduling. Since the target is a small integer (maximum schedule length $16$), the base model quickly saturates to Pass@k $\approx 1.0$ as $k$ increases. While the RLVR-trained policy attains slightly lower Pass@k  at large $k$, it achieves much higher self-consistency (about $0.68$ vs. $\sim\!0.24$ at k=256k{=}256), indicating more stable predictions across samples. Under exact sequence ID matching, RLVR substantially outperforms the base model: at k=256k{=}256, RLVR reaches $\mathrm{Pass@k}\approx 0.64$ compared to $0.14$, with $\mathrm{SC}\approx 0.34$ compared to $0.004$, reflecting a clear improvement in sequence-level fidelity. These results support prior findings that answer-only Pass@k can overstate model capability [17], whereas $\mathrm{Acc}_{\text{ids}}$ and SC provide more faithful measures of reasoning quality. RLVR improves both by reinforcing verified reasoning trajectories [17].
Figure 2: Performance comparison of Base, RL($r_{\text{ans}}$), and RL($r_{\text{ans+fmt}}$) with Qwen2.5-7B.
@@
Figure 3: Performance comparison of RL models trained with $(r_{ids,exa})$ and $(r_{ids,pre})$ models on the Activity and LIS task with the Qwen2.5-7B model.
@@ 3.2 Training with Sequence Rewards
3.3 Sorting Performance and Reward on Activity Task
@@
Figure 4 summarizes sorting quality for the four models. Both sequence-reward policies ($r_{\text{ids,exa}}$ and $r_{\text{ids,pre}}$) improve exact sorting relative to the base and $r_{\text{ans}}$ models (from $0.1\%/0.4\%$ to $2.0\%/1.9\%$) and increase the mean LCS (from $0.238/0.290$ to $0.517/0.444$). However, absolute exact-sorting accuracy remains very low ($\approx\!2\%$), even though these same policies achieve high sequence correctness at evaluation (e.g., $\mathrm{Pass@256}_{\text{ids}}\!\approx\!0.60$ on Activity). This gap indicates that the sorting step is not reliably driving the final schedule, despite the frequent appearance of a “sorted” preface.
Figure 4: Sorting acc. on Activity Scheduling
Training with Sorting-Match Reward. We tested a sorting-match reward $r_{\text{sort}}$ to encourage the first step of activity scheduling, but this caused collapse: both $\mathrm{Acc}_{\text{ans}}$ and $\mathrm{Acc}_{\text{ids}}$ dropped to nearly $0\%$ as the model output only the sorted sequence without enforcing non-overlap. Adding $r_{\text{sort}}$ to $r_{\text{ans}}$ and $r_{\text{ids}}$ restored performance to the $r_{\text{ids}}$ baseline but yielded no further gains. Curriculum experiments further showed that extended pretraining on $r_{\text{sort}}$ hindered recovery under $r_{\text{ans}}$, with longer exposure preventing learning even on the training set (see Appendix H).
Takeaway 3 In Activity Scheduling, RLVR with sequence rewards improves answer and sequence accuracy, but sorting accuracy remains very low. Models often emit a superficial “sorted” preface that neither matches the canonical order nor drives the final schedule, highlighting a disconnect between surface outputs and the underlying decision rule [13].
@@ 4 Limitations & Future Work
Our study investigates two verifiable reasoning tasks (Activity Scheduling and LIS) with a single base model (Qwen2.5-7B-Instruct). While this controlled setup helps isolate RLVR dynamics, conclusions may not generalize across tasks, model families, or scales, as prior work has shown diverse shortcut behaviors and model-dependent failure modes [8]. To partially address this, we provide additional results for Llama-3.1-8B in Appendix F.
Future work could expand to a broader range of models and problem domains, and employ mechanistic interpretability to probe the circuits and dynamics underlying RLVR-induced behaviors [10]. Such analysis may clarify whether improvements reflect genuine task learning or superficial strategies that exploit evaluation metrics. Our findings highlight this tension, emphasizing the need for careful evaluation and diagnostics when claiming improvements in reasoning from RLVR.
@@ Appendix D LIS Task Response Length and Entropy
Figure 7 further compares entropy during training. Entropy drops sharply for $r_{\text{ans}}$, mirroring the collapse in response length, whereas $r_{\text{ans+fmt}}$ maintains higher entropy throughout training. Despite these qualitative differences, both RLVR-trained policies converge to similar levels of $\mathrm{Acc}_{\text{ans}}$ and $\mathrm{Acc}_{\text{ids}}$, suggesting reliance on comparable heuristics expressed through different generation styles. For example, with $r_{\text{ans+fmt}}$ nearly all responses contain Python code (100%, compared to 35.1% for the base model), yet the model does not execute the code step by step to derive the answer. This observation echoes findings by Shao et al. [8], who show that RLVR can amplify spurious reward signals in the Qwen family, encouraging models to emit structured but ultimately superficial outputs.
@@ Appendix F Llama Model Performance
G.1 Heuristics Analysis for LIS
Model $R^{2}_{\text{test}}$ $\text{MAE}_{\text{test}}$
Base -0.002 2.745
$r_{\text{ans}}$ 0.741 0.269
$r_{\text{ids,exa}}$ 0.781 0.289
$r_{\text{ids,pre}}$ 0.841 0.227
$r_{\text{ans+fmt}}$ 0.867 0.209
Table 1: Predictive fit of LIS
features on model outputs.
Appendix H Training with Sorting-Match Reward.
We investigated whether directly rewarding the sorting step, the first stage of the greedy activity-scheduling algorithm, could yield additional benefits by training with a sorting-match reward $r_{\text{sort}}$. This approach backfired: both $\mathrm{Acc}_{\text{ans}}$ and $\mathrm{Acc}_{\text{ids}}$ collapsed to nearly $0\%$, as the model simply output the sorted sequence without enforcing the non-overlap constraint. A combined objective with equal weights on $r_{\text{ans}}$, $r_{\text{ids}}$, and $r_{\text{sort}}$ restored performance to the level of $r_{\text{ids}}$ alone but did not improve sorting on the test set, suggesting that sorting in isolation provides no complementary benefit. Curriculum experiments confirmed this: pretraining with $r_{\text{sort}}$ for 10 PPO steps allowed recovery under $r_{\text{ans}}$, but 20 or 30 steps severely hindered it, with 30-step pretraining preventing any improvement even on the training set (Figure 12).

### v2

@@ 1 Introduction
Large language models (LLMs) have recently advanced rapidly on mathematical and programming benchmarks [guo2025deepseek, jaech2024openai, wu2024reasoning, comanici2025gemini, yang2025qwen3]. A key driver is Reinforcement Learning with Verifiable Rewards (RLVR), which fine-tunes pretrained models against automatically checkable signals such as exact answers or unit tests [guo2025deepseek, yu2025dapo]. This paradigm eliminates reliance on human annotation, enabling scalable training on large problem sets and delivering consistent gains on challenging reasoning tasks [el2025competitive].
Despite strong empirical gains, the nature of RLVR’s improvements remains unclear. Studies show it often boosts accuracy while reducing exploration [wu2025invisible, cui2025entropy, yue2025does, wang2025beyond], with base model ability acting as a ceiling [yue2025does, wu2025invisible]. Gains largely reflect re-weighting existing solutions, and reasoning coverage can even contract at larger sample sizes. Standard $Pass@K$ further over-credits “lucky” completions [wen2025reinforcement], whereas stricter metrics like CoTCoT-$Pass@K$ are more reliable but harder to scale. Reward design can also yield surprising effects: a single example can rival large-scale training [wang2025reinforcement], and even spurious rewards can drive improvements in models with strong procedural biases [shao2025spurious]. Overall, RLVR appears to stabilize existing competencies rather than induce new reasoning strategies.
@@ 2.1 Tasks
Activity Scheduling. Each activity $i$ has start and finish times $(s_{i},f_{i})$ with $s_{i}<f_{i}$, and intervals are half-open $s_{i},f_{i})$. The goal is to select a maximum subset of non-overlapping activities [wikipedia_activity_selection]. Instances are constructed so that the greedy earliest-finish-time algorithm with deterministic tie-breaking yields the unique optimum, reported as IDs sorted by $f_{i}$ (ties by smaller $i$).
Longest Increasing Subsequence (LIS). Given $a_{1},\dots,a_{n}\in\mathbb{Z}$, find indices $1\leq i_{1}<\cdots<i_{k}\leq n$ maximizing $k$ with $a_{i_{1}}<\cdots<a_{i_{k}}$. Uniqueness is enforced via an $O(n^{2})$ dynamic-programming count, and the LIS is reconstructed with patience sorting in $O(n\log n)$ [wikipedia_lis].
@@ 2.2 Reward Functions
$r_{\text{ids,exa}}(y)\;=\;\mathbb{I}\!\bigl[\hat{s}(y)=s^{\star}\bigr].$
@@
$r_{\text{ids,pre}}(y)\;=\;\max\Bigl\{\,0,\ \frac{m(y)}{L}\;-\;\gamma\,\mathbb{I}\!\bigl[\,\hat{s}(y)=\varnothing\ \lor\ |\hat{s}(y)|\neq L\,\bigr]\Bigr\}.$
@@ 2.3 Training & Evaluation
We fine-tune Qwen2.5-7B-Instruct [qwen2.5] with GRPO [shao2024deepseekmath, guo2025deepseek] using the verl framework [sheng2024hybridflow]. Unless noted, max generation length is $T_{\max}=2048$ (extended to $7680$ for LIS with $r_{\text{ans+fmt}}$). Each PPO update uses $256$ prompts with $8$ rollouts from vLLM [kwon2023efficient], trained for $20$ epochs ($120$ updates) at learning rate $10^{-6}$ and no KL penalty ($\beta_{\text{KL}}=0$). Training prompts match evaluation format.
@@
  • Pass@k [yue2025does]. Success under a $k$-sample budget, defined by whether at least one of the $k$ generations is correct:
@@
  • Self-consistency (SC) [wang2022self]. Agreement between the majority prediction aggregated over $k$ generations and the ground truth:
@@ 3.1 Training with Exact Answer Reward
Activity Scheduling. Since the target is a small integer (maximum schedule length $16$), the base model quickly saturates to Pass@k $\approx 1.0$ as $k$ increases. While the RLVR-trained policy attains slightly lower Pass@k  at large $k$, it achieves much higher self-consistency (about $0.68$ vs. $\sim\!0.24$ at k=256k{=}256), indicating more stable predictions across samples. Under exact sequence ID matching, RLVR substantially outperforms the base model: at k=256k{=}256, RLVR reaches $\mathrm{Pass@k}\approx 0.64$ compared to $0.14$, with $\mathrm{SC}\approx 0.34$ compared to $0.004$, reflecting a clear improvement in sequence-level fidelity. These results support prior findings that answer-only Pass@k can overstate model capability [wen2025reinforcement], whereas $\mathrm{Acc}_{\text{ids}}$ and SC provide more faithful measures of reasoning quality. RLVR improves both by reinforcing verified reasoning trajectories [wen2025reinforcement].
Figure 2: Performance comparison of Base, RL($r_{\text{ans}}$), and RL($r_{\text{ans+fmt}}$) models on the Activity and LIS task with the Qwen2.5-7B model.
@@
Figure 3: Performance comparison of RL models trained with $r_{ids,exa}$ and $r_{ids,pre}$.
@@ 3.2 Training with Sequence Rewards
3.3 Sorting Performance on Activity Task
@@
Figure 4 summarizes sorting quality for the four models. Both sequence-reward policies ($r_{\text{ids,exa}}$ and $r_{\text{ids,pre}}$) improve exact sorting relative to the base and $r_{\text{ans}}$ models (from $0.17\%/0.43\%$ to $2.01\%/1.85\%$) and increase the mean LCS (from $0.248/0.290$ to $0.517/0.444$). However, absolute exact-sorting accuracy remains very low ($\approx\!2\%$), even though these same policies achieve high sequence correctness at evaluation (e.g., $\mathrm{Pass@256}_{\text{ids}}\!\approx\!0.60$ on Activity). This gap indicates that the sorting step is not reliably driving the final schedule, despite the frequent appearance of a “sorted” preface.
Figure 4: Sorting accuracy and LCS across models.
3.4 Training with Sorting-Match Reward
In §3.2, we observed performance gains in $\mathrm{Acc}_{\text{ans}}$ when training with sequence-aware rewards. This raises the question of whether rewarding the sorting step, which is the first stage of the greedy activity-scheduling algorithm, could provide additional benefit. To test this, we trained a model with the sorting-match reward $r_{\text{sort}}$.
Surprisingly, this led to a catastrophic collapse: both $\mathrm{Acc}_{\text{ans}}$ and $\mathrm{Acc}_{\text{ids}}$ fell to nearly $0\%$. Inspection of the outputs revealed that the model consistently produced the sorted sequence itself as the final answer, without applying the non-overlap constraint required for valid schedules.
We next trained a model using a combined objective with equal weights on $r_{\text{ans}}$, $r_{\text{ids}}$, and $r_{\text{sort}}$. This configuration restored $\mathrm{Acc}_{\text{ans}}$ and $\mathrm{Acc}_{\text{ids}}$ to levels comparable to the $r_{\text{ids}}$ model, but it did not improve sorting performance on the test set. Thus, learning to sort in isolation does not confer complementary benefits for solving the activity-scheduling task.
To probe further, we ran a curriculum experiment: training with $r_{\text{sort}}$ alone for the first 10, 20, or 30 PPO updates, and then switching to $r_{\text{ans}}$ for the remainder of the 120 training steps. As shown in Figure 5, models trained with $r_{\text{sort}}$ for only 10 steps were able to recover under $r_{\text{ans}}$, but longer exposure (20 or 30 steps) severely hindered recovery. In fact, after 30 steps of $r_{\text{sort}}$, the model failed to improve $\mathrm{Acc}_{\text{ans}}$ even on the training data.
Takeaway 3 In Activity Scheduling, RLVR with sequence rewards improves answer and sequence accuracy, but sorting accuracy remains very low. Models often emit a superficial “sorted” preface that neither matches the canonical order nor drives the final schedule, highlighting a disconnect between surface outputs and the underlying decision rule [turpin2023language].
3.5 Heuristics Analysis for LIS
Model $R^{2}_{\text{test}}$ $\text{MAE}_{\text{test}}$
Base -0.002 2.745
$r_{\text{ans}}$ 0.741 0.269
$r_{\text{ids,exa}}$ 0.781 0.289
$r_{\text{ids,pre}}$ 0.841 0.227
$r_{\text{ans+fmt}}$ 0.867 0.209
Table 1: Predictive fit of LIS
features on model outputs.
@@ 4 Limitations & Future Work
Our study investigates two verifiable reasoning tasks (Activity Scheduling and LIS) with a single base model (Qwen2.5-7B-Instruct). While this controlled setup helps isolate RLVR dynamics, conclusions may not generalize across tasks, model families, or scales, as prior work has shown diverse shortcut behaviors and model-dependent failure modes [shao2025spurious]. To partially address this, we provide additional results for Llama-3.1-8B in Appendix F.
Future work could expand to a broader range of models and problem domains, and employ mechanistic interpretability to probe the circuits and dynamics underlying RLVR-induced behaviors [sharkey2025open]. Such analysis may clarify whether improvements reflect genuine task learning or superficial strategies that exploit evaluation metrics. Our findings highlight this tension, emphasizing the need for careful evaluation and diagnostics when claiming improvements in reasoning from RLVR.
@@ Appendix D LIS Task Response Length and Entropy
Figure 8 further compares entropy during training. Entropy drops sharply for $r_{\text{ans}}$, mirroring the collapse in response length, whereas $r_{\text{ans+fmt}}$ maintains higher entropy throughout training. Despite these qualitative differences, both RLVR-trained policies converge to similar levels of $\mathrm{Acc}_{\text{ans}}$ and $\mathrm{Acc}_{\text{ids}}$, suggesting reliance on comparable heuristics expressed through different generation styles. For example, with $r_{\text{ans+fmt}}$ nearly all responses contain Python code (100%, compared to 35.1% for the base model), yet the model does not execute the code step by step to derive the answer. This observation echoes findings by shao2025spurious, who show that RLVR can amplify spurious reward signals in the Qwen family, encouraging models to emit structured but ultimately superficial outputs.